OpenStudy (anonymous):

Choose the system of equations which matches the following graph. MULTIPLE CHOICE

OpenStudy (anonymous):

3x – 6y = 12 x – 2y = 10 3x + 6y = 12 x + 2y = 10 3x – 6y = 12 x + 2y = 10 3x + 6y = 12 x – 2y = 10

OpenStudy (anonymous):

....see the x and y intercepts in the graph..watch for the set of equations which satisfy their respective intercepts.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do you want the answer or you can do it?

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

No..watch the y intercept of the second equation..its (0,-5)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks i appreciate it!

OpenStudy (anonymous):

welcome:)